Blayton40:

150x + 80y + 12z = 2,466 380x + 95y + 48z = 5,419 210x + 130y + 60z = 3,960 220x + 225y = 4,445 540x + 270y = 8,370 → y = −2x + 31 x = 11 Finish solving the system using. What are y and z?

darkknight:

It seems that you have a few different questions on this post. Can you clarify what the question is? Is it 150x + 80y + 12z = 2,466 380x + 95y + 48z = 5,419 210x + 130y + 60z = 3,960 or...